# Open-vocabulary Multimodal Emotion Recognition: Dataset, Metric, and Benchmark

### v1

@@ Abstract
Multimodal Emotion Recognition (MER) is an important research topic. This paper advocates for a transformative paradigm in MER. The rationale behind our work is that current approaches often rely on a limited set of basic emotion labels, which do not adequately represent the rich spectrum of human emotions. These traditional and overly simplistic emotion categories fail to capture the inherent complexity and subtlety of human emotional experiences, leading to limited generalizability and practicality. Therefore, we propose a new MER paradigm called Open-vocabulary MER (OV-MER), which encompasses a broader range of emotion labels to reflect the richness of human emotions. This paradigm relaxes the label space, allowing for the prediction of arbitrary numbers and categories of emotions. To support this transition, we provide a comprehensive solution that includes a newly constructed database based on LLM and human collaborative annotations, along with corresponding metrics and a series of benchmarks. We hope this work advances emotion recognition from basic emotions to more nuanced emotions, contributing to the development of emotional AI.
@@ 1 Introduction
Research on emotions has a history spanning two centuries. As early as the 19th century, Charles Darwin conducted pioneering research about the evolutionary origins and possible purposes of emotions, explaining the emotional expressions of humans and animals [1]. In 1884, William James revealed the process of emotion generation, noting that stimuli trigger activities in the autonomic nervous system, which in turn produces an emotional experience in the brain [2]. With the rapid development of AI, emotions have garnered increasing attention. [3] highlighted the importance of emotions: The question is not whether intelligent machines can have any emotions, but whether machines can be intelligent without any emotions.
Humans convey emotions through various modalities, which gives rise to the task of Multimodal Emotion Recognition (MER) [4]. Its main goal is to integrate multimodal clues to identify human emotional states. The first step in recognizing emotions is how to model them. Currently, emotion models can be broadly classified into two categories: dimensional models and discrete models. The former uses multi-dimensional space to describe emotions, with the arousal-valence dimension being widely adopted; valence describes the pleasantness of a stimulus, while arousal indicates the intensity of emotion provoked by a stimulus [5]. However, these definitions require specialized psychological knowledge, making them abstract and difficult for the ordinary person to understand. This can lead to discrepancies between different annotators, posing challenges for subsequent applications.
Discrete models align more closely with human understanding of emotions. [6] proposed the basic emotion theory, suggesting that there are six basic emotions: anger, disgust, fear, happiness, sadness, and surprise. This theory is widely used in MER, where researchers typically limit the label space to these basic emotions and use multiple annotators to select the most likely label through majority voting. We refer to this task as One-hot MER (OH-MER). Considering that emotions can be compound, researchers further propose Multi-label MER (ML-MER), allowing each sample to have multiple labels [7]. However, both OH-MER and ML-MER generally have limited label spaces and numbers. Nevertheless, [8] pointed out that humans can express approximately 34,000 different emotions. Therefore, restricting the label space and number will overlook some nuanced emotions.
@@
Figure 1: Comparison. (a) Task Comparison: We compare the differences among three tasks (one-hot MER, multi-label MER, and open-vocabulary MER) across three aspects (label space, label number, and annotation manner). (b) Label Comparison: We provide an example to visualize the one-hot and OV labels. This paper uses emotion-related descriptions as a bridge for OV label extraction. We observe that OV labels contain richer emotions. See Appendix B for more examples.
@@
In this paper, we extend traditional MER to Open-vocabulary MER (OV-MER), where we allow the prediction of emotions across any number and categories. Figure 1(a) provides comparisons between different tasks. To facilitate further research, we build an initial dataset, define evaluation metrics, and develop baselines: (1) For the dataset, we propose a human-LLM collaboration strategy. Compared to human-only annotation, our strategy can leverage LLM to enhance the label richness (see Section 2); (2) For evaluation metrics, since there is no fixed label space, the model may predict closely related but differently expressed emotions. For example, the ground truth is joyful and the model predicts happy. To provide more reliable evaluation results, we first group similar emotions and specifically design metrics for this task (see Section 3); (3) For baselines, traditional classifiers rely on fixed label spaces. However, OV-MER does not restrict the label space or number of labels, necessitating the definition of new solutions (see Section 4).
@@
A natural question arises: why is open vocabulary so important for MER? The reason is that OV can generate richer and more nuanced emotional representations, leading to more accurate and reliable MER. As illustrated in Figure 1(b), labeling an emotion solely as surprise is not sufficiently informative—we cannot ascertain whether the surprise is positive or negative, which limits its practicality, such as in human-computer interaction (HCI) applications. In contrast, our OV-MER provides emotions like surprise along with additional descriptors such as nervous and dissatisfied, offering a more comprehensive and insightful understanding of the emotional state. Therefore, OV-MER facilitates the transition from basic to nuanced emotion recognition, advancing the development of emotion AI. In summary, we make the following key contributions:
@@
  • Paradigm. We propose a new paradigm in MER called OV-MER. This paradigm transitions from traditional MER to a framework that enables the prediction of any number and category of emotions, thereby advancing emotion AI toward real-world applicability by capturing the full spectrum of human emotions.
@@
  • Groundwork. We lay the groundwork for OV-MER by constructing datasets, defining evaluation metrics, and proposing effective solutions. Our dataset enhances label richness through a human-LLM collaboration strategy. Meanwhile, we introduce new evaluation metrics that leverage emotional relevance to achieve more reliable results.
@@
  • Benchmark. We establish benchmarks for OV-MER by conducting extensive experiments on various baselines and providing detailed analysis. This task can serve as an important evaluation benchmark for MLLMs, challenging their ability to integrate multimodal clues and capture subtle temporal variations in emotional expression.
@@ 2 The OV-MERD Dataset Construction
Although the concept of OV-MER is intuitive and holds great promise, its practical implementation faces significant challenges. The main difficulty lies in the broad and subtle range of human emotions, making comprehensive labeling a complex task. Traditional annotation methods are limited by their predefined emotion categories, which are often insufficient for the needs of OV-MER. Therefore, as shown in Figure 2, we propose a human-LLM collaboration strategy that consists of two steps: CLUE-Multi generation and emotion label extraction. Ultimately, we create a dataset called OV-MERD, which offers a richer set of emotions compared to existing datasets (see Table 1).
2.1 CLUE-Multi Generation
During the annotation process, we observe that descriptions generated through human-LLM collaboration are more detailed. This indicates that when annotations are solely performed by humans, they tend to focus on major clues while neglecting minor ones (see Section 5 for more experimental results). In this section, we provide a detailed overview of our human-LLM collaboration strategy.
Figure 2: Dataset construction. (a) CLUE-Multi Generation: For audio and video, we use ALLM and VLLM to extract initial clues, followed by two rounds of manual checks to eliminate errors and duplicates while adding missing content. Each round involves multiple annotators, with no overlap between annotators in the two rounds. Finally, we merge the checked clues with text to generate CLUE-Multi. (b) Ground-truth OV Label Extraction: There are certain differences in the labels extracted from different languages. To eliminate language influence and achieve consensus labels, we merge these labels and conduct manual checks.
Pre-annotation.
Initially, we attempt to directly annotate visual and acoustic clues. However, the descriptions obtained in this way cannot cover all information. Therefore, we explore using other models for pre-annotation. (1) For video, given the strong visual understanding capabilities of GPT-4V (“gpt-4-vision-preview”), we use it as video LLM (VLLM) to pre-label visual clues. Since GPT-4V only supports image input, we uniformly sample each video and convert it into an image sequence. Considering the relatively short duration of our video, we sample three frames as input. (2) For audio, we use the open-source SALMONN [9] as audio LLM (ALLM) for pre-annotation, because GPT-4V does not support audio input.
Manual Check.
As part of our quality assurance procedures, we perform a detailed examination of the pre-annotated results. For visual clues, GPT-4V may generate hallucinated responses, i.e., clues that do not actually exist. Additionally, there are repeated expressions and some temporal association clues are missing. Therefore, we hire annotators to eliminate errors and duplicates, as well as add missing information. For acoustic clues, ALLM struggles to capture emotion-related acoustic features. Hence, we hire annotators to focus on the speaker’s intonation and other emotional clues. To reduce subjective bias, we conduct two rounds of manual checks, each involving different annotators. These annotators are experts in affective computing and are familiar with the definitions of emotions. Ultimately, these checked clues can accurately reflect the video content. Appendix G provides the annotation guideline and layout of the annotation platform.
CLUE-Multi Generation.
Subsequently, we leverage the semantic reasoning capabilities of LLM to integrate all clues. Specifically, we use GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as the LLM and request it to merge textual, audio, and visual clues to infer emotional states. The output of this process is an emotion-related description, denoted as CLUE-Multi (see Figure 2). Overall, this process reflects the collaboration between humans and LLMs. In Section 5, we compare human-only annotation with this strategy, observing that the latter produces more comprehensive descriptions. In Appendix I, we conduct a further in-depth analysis of CLUE-Multi, demonstrating that it provides a complete representation of emotions, supporting both discrete and dimensional emotion recognition.
2.2 Ground-truth OV Label Extraction
Label Extraction.
After that, we use the LLM to extract emotion labels from CLUE-Multi. This process relies on GPT-3.5, which we request to identify emotional states based on the provided descriptions without restricting the label space. See Appendix D for more details.
Label Filtering.
We further explore the language impact. In Figure 2, we first extract emotion labels from English and Chinese descriptions, resulting in $Y_{EE}$ and $Y_{CC}$. Then, we translate these labels into the other language, yielding $Y_{EC}$ and $Y_{CE}$. After that, we calculate their similarity to assess language influence. In this section, we rely on the Jaccard similarity coefficient to measure the similarity between two sets (see Appendix H). Experimental results indicate that labels extracted from different languages exhibit some differences. For example, the similarity score between $Y_{EE}$ and $Y_{CE}$ is 0.82, which may be due to the varying definitions of emotions in different languages. To eliminate language influence and achieve consensus labels, we merge the labels extracted from both languages and conduct manual checks. These checked labels are regarded as the ground truth.
Table 1: Dataset comparison. “A”, “V”, and “T” are abbreviations for audio, video, and text, respectively. See Appendix F for a comprehensive comparison with existing datasets.
@@ 2.2 Ground-truth OV Label Extraction
OV-MERD (Ours) A,V,T Discrete Emotion 248
@@ 2.3 OV-MERD Dataset
Finally, we construct a dataset called OV-MERD. This dataset is an extension of MER2023 [20], from which we evenly select samples with further annotations. Table 1 compares OV-MERD with existing datasets. We observe that our OV-MERD dataset contains 248 emotion categories and most samples have 2 to 4 labels, far exceeding those in current datasets. In Appendix E, we observe that OV-MERD encompasses a broader range of emotions, including some that have been rarely discussed in previous research, such as shy, nervous, and grateful.
@@ 3 Evaluation Metric
Defining evaluation metrics for OV-MER presents significant challenges: (1) OV-MER supports predicting emotions of any category. Thus, the model may predict closely related but differently expressed emotions. To provide more reliable evaluation results, we first group the emotions based on their similarities. (2) OV-MER allows for the prediction of an arbitrary number of labels. Thus, traditional evaluation metrics designed for a fixed number of labels may not be applicable. In this section, we propose set-based evaluation metrics specifically tailored for OV-MER.
3.1 Grouping
We propose two grouping strategies: one based on GPT and the other based on the emotion wheel (EW) [22]. In the following experiments, we default to using GPT-based grouping.
GPT-based Grouping.
The most direct approach is to use GPT-3.5 to group all labels based on their similarity: Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing emotions with the same meaning. Directly output the results. The output format should be a list containing multiple lists. However, the evaluation results may be affected by the API version. For example, if OpenAI deprecates an old API, the results based on that API will become difficult to reproduce. Additionally, this process is costly; evaluating our OV-MERD dataset costs about \$1, and costs will further increase if the dataset is expanded in the future. Therefore, we attempt to find a replacement for GPT-based grouping.
Figure 3: Emotion wheel W1.
EW-based Grouping.
[…]
Take W1 as an example (see Figure 3). Before calculating the metrics, we define some symbols. We group the labels by their levels from the innermost to the outermost as $L_{w_{1}}^{1}$, $L_{w_{1}}^{2}$, and $L_{w_{1}}^{3}$. Specifically, $L_{w_{1}}^{1}=\{\text{mad},\cdots,\text{scared}\}$. Next, we define a function $m_{w_{1}}^{i\rightarrow j}(\cdot)$ that maps the labels in $L_{w_{1}}^{i}$ to the corresponding labels in $L_{w_{1}}^{j}$. From inner to outer ($i<j$), $m_{w_{1}}^{i\rightarrow j}(\cdot)$ is a many-to-one mapping; from outer to inner ($i>j$), $m_{w_{1}}^{i\rightarrow j}(\cdot)$ is a one-to-many mapping. For example, $m_{w_{1}}^{2\rightarrow 1}(\mbox{angry})=\{\mbox{mad}\}$ and $m_{w_{1}}^{1\rightarrow 2}(\mbox{mad})=\{\mbox{offended},\cdots,\mbox{humiliated}\}$. We collect all the labels from these emotion wheels and represent them as EW, i.e., $\{L_{w_{i}}^{j},1\leq i\leq 5,1\leq j\leq 3\}$. We denote the labels in EW as $y_{w}$.
[…]
(1) M1. We use $g^{\prime}(\cdot)$ to map all labels to their corresponding $y_{f}$.
(2) M2. We use $f^{\prime}(g^{\prime}(\cdot))$ to map all labels to their corresponding $y_{w}$.
(3) M3. We use the emotion wheel during metric calculation. Specifically, we first use $f^{\prime}(g^{\prime}(\cdot))$ to map all labels to their corresponding $y_{w}$. Then, we define two grouping functions, L1 and L2. For L1, we map all labels to their corresponding $L_{w_{i}}^{1}$:
For L2, we map all labels to their corresponding $L_{w_{i}}^{2}$:
3.2 Metric Definition
Then, we convert the above emotion grouping information into a function $G(\cdot)$, which can map each label to its group ID. Specifically, suppose $\{y_{i}\}_{i=1}^{M}$ and $\{\hat{y}_{i}\}_{i=1}^{N}$ are the ground truth and predictions, where $M$ and $N$ are the number of labels. We first map each label into its group ID:
$\mathcal{Y}=\{G(x)|x\in\{y_{i}\}_{i=1}^{M}\},\hat{\mathcal{Y}}=\{G(x)|x\in\{\hat{y}_{i}\}_{i=1}^{N}\}.$ (3)
Finally, we design set-based metrics for performance evaluation. Specifically, $\mbox{Precision}_{\mbox{s}}$ indicates the number of correctly predicted labels; $\mbox{Recall}_{\mbox{s}}$ indicates whether the prediction covers all ground truth; and $\mbox{F}_{\mbox{s}}$ is the harmonic mean of two metrics, which is used for the final ranking:
4 Baselines for OV-MER
4.1 CLUE Generation
Figure 2 illustrates the generation process of CLUE-Multi, where we combine text with manually checked visual and acoustic clues. In this section, we further introduce the following variants.
CLUE-A/T/V.
To reveal the modality impact, we propose three variants of CLUE-Multi: CLUE-Audio, CLUE-Text, and CLUE-Video. Their generation process is illustrated in Figure 4. Specifically, since audio contains both acoustic information (e.g., the speaker’s intonation) and content, we merge the checked acoustic clues and text to generate CLUE-Audio; for CLUE-Text, we only use the text to infer emotional states; since the visual content does not contain audio and text, for CLUE-Video, we only use the checked visual clues. See Appendix J for more examples.
CLUE-MLLM.
MLLMs can address various multimodal tasks, with the main idea being to align information from other modalities with LLMs. Since emotion recognition relies on temporal information, we choose models that support at least video or audio. To generate CLUE-MLLM, we first use ALLM or VLLM to extract emotion-related descriptions, and then combine these descriptions with text using LLM. Compared with CLUE-Multi, this process does not use manually checked clues. Appendix K provides the model card and relevant prompts. All models are implemented in PyTorch, and all inference processes are carried out using a 32G NVIDIA Tesla V100 GPU.

### v2

# OV-MER: Towards Open-Vocabulary Multimodal Emotion Recognition

@@ Abstract
Multimodal Emotion Recognition (MER) is a critical research area that seeks to decode human emotions from diverse data modalities. However, existing machine learning methods predominantly rely on predefined emotion taxonomies, which fail to capture the inherent complexity, subtlety, and multi-appraisal nature of human emotional experiences, as demonstrated by studies in psychology and cognitive science. To overcome this limitation, we advocate for introducing the concept of open vocabulary into MER. This paradigm shift aims to enable models to predict emotions beyond a fixed label space, accommodating a flexible set of categories to better reflect the nuanced spectrum of human emotions. To achieve this, we propose a novel paradigm: Open-Vocabulary MER (OV-MER), which enables emotion prediction without being confined to predefined spaces. However, constructing a dataset that encompasses the full range of emotions for OV-MER is practically infeasible; hence, we present a comprehensive solution including a newly curated database, novel evaluation metrics, and a preliminary benchmark. By advancing MER from basic emotions to more nuanced and diverse emotional states, we hope this work can inspire the next generation of MER, enhancing its generalizability and applicability in real-world scenarios.
@@ 1 Introduction
Research on emotions has a history spanning two centuries. As early as the 19th century, Charles Darwin conducted pioneering research about the evolutionary origins and possible purposes of emotions, explaining the emotional expressions of humans and animals [1]. In 1884, James revealed the process of emotion generation, noting that stimuli trigger activities in the autonomic nervous system, which in turn produces an emotional experience in the brain [2]. With the rapid development of AI, emotions have garnered increasing attention [3].
The basis of Multimodal Emotion Recognition (MER) lies in the effective modeling of emotions. Current emotion models are primarily categorized into two types: dimensional and discrete models. Dimensional models, particularly those based on psychological theories such as the Circumplex Model of Affect [4], represent emotions within a continuous, multi-dimensional space. The most widely adopted framework uses two or three primary dimensions: valence (the pleasantness-unpleasantness continuum), arousal (the activation-deactivation level), and dominance (the degree of control perceived). These dimensions allow for the quantification of emotional states into measurable numerical values [5]. However, this sophisticated numerical representation demands specialized psychological expertise for accurate interpretation, making it abstract and less descriptive to the general public. This abstraction can result in inconsistencies among different annotators, particularly in complex emotional states that fall between the primary dimensions, thereby complicating subsequent applications and potentially affecting the reliability of emotion recognition systems.
Discrete models, which categorize emotions into distinct classes, tend to mirror the way people naturally perceive and express emotions in daily life. [6] proposed the basic emotion theory, suggesting that there are six basic emotions: anger, disgust, fear, happiness, sadness, and surprise. This theory is widely used in MER, where researchers typically limit the label space to these basic emotions and use multiple annotators to select the most likely label through majority voting. We refer to this task as One-hot MER (OH-MER). Considering that emotions can be compound, researchers further propose Multi-label MER (ML-MER), allowing each sample to have multiple labels [7]. However, both OH-MER and ML-MER generally have limited label spaces. [8] pointed out that humans can express approximately 34,000 distinct emotions. Although some efforts have been made to expand label spaces with more emotional categories, current approaches still fail to capture emotion diversity, inevitably overlooking some of these nuanced emotions.
@@
Figure 1: Comparison. (a) Task Comparison: We compare the differences among three tasks (one-hot MER, multi-label MER, and OV-MER) across three aspects (label space, label number, and annotation manner). An in-depth comparison is provided in the Appendix A; (b) Label Comparison: We provide an example to visualize the one-hot and OV labels. More examples are provided in Appendix E. Since the original video contains real people, we use DemoAI to remove personal information to address copyright concerns. In this paper, we use emotion-related descriptions as a bridge to extract OV labels. We observe that OV labels offer a more insightful understanding of the emotional state.
Figure 2: Dataset construction. (a) CLUE-Multi Generation: For audio and video, we use audio LLM (ALLM) and video LLM (VLLM) to extract initial clues, followed by two rounds of manual checks to eliminate errors and duplicates while adding missing content. Each round involves multiple annotators, with no overlap between annotators in the two rounds. Finally, we merge the checked clues with text to generate CLUE-Multi. (b) Ground-truth OV Label Extraction: There are certain differences in the labels extracted from different languages. To eliminate language influence and achieve consensus labels, we merge these labels and conduct manual checks. These checked labels are regarded as the ground truth.
@@
In this paper, we introduce a new MER paradigm, open-vocabulary MER (OV-MER), by introducing the concept of open-vocabulary into MER, enabling the prediction of arbitrary emotion categories. Figure 1(a) provides a comparison between different tasks, and an in-depth comparison is provided in Appendix A. To support this shift, we build a dataset, define evaluation metrics, and develop solutions. (1) Dataset: we propose a human-LLM collaboration strategy to construct the dataset. Compared to human-only annotation, our strategy can leverage LLM to enhance the label richness; (2) Metrics: since there is no fixed label space, the model may predict closely related but differently expressed emotions (e.g., joyful and happy). To provide more reliable evaluation results, we first group similar emotions and specifically design metrics for this task; (3) Solutions: traditional classifiers rely on fixed label spaces. However, OV-MER does not restrict the label space, necessitating the definition of new solutions.
@@
A natural question arises: why is OV-MER so important? A simple answer is that it naturally aligns with the way emotions are expressed in our real-life interactions, leading to more accurate and human-centered MER. As illustrated in Figure 1(b), labeling an emotion solely as happy is not sufficiently informative. In contrast, OV-MER provides emotions like mockery, offering a more comprehensive and insightful understanding of the emotional state. Therefore, OV-MER facilitates the transition from basic to nuanced emotion recognition, advancing the development of emotion AI. Appendix C provides more detailed motivation. In summary, we make the following key contributions:
@@
  • Paradigm. We propose a new paradigm in MER, called OV-MER. This paradigm transitions from traditional MER to a framework that enables the prediction of any number and category of emotions, thereby advancing emotion AI toward real-world applicability by capturing the full spectrum of human emotions.
@@
  • Groundwork. We lay the groundwork for OV-MER by constructing datasets, defining evaluation metrics, and proposing solutions. Our dataset enhances label richness through human-LLM collaboration. Meanwhile, we introduce new evaluation metrics that leverage emotional relevance to achieve more reliable results.
@@
  • Benchmark. We build zero-shot benchmarks for OV-MER through extensive experiments and detailed analysis. This task can serve as an important evaluation benchmark for multimodal LLMs (MLLMs), challenging their ability to integrate multimodal clues and capture subtle temporal variations in emotional expression.
  • Experiments. Our intensive experimental results not only demonstrate the strength of our methods but also prove that OV-MER can effectively enhance the presentation ability of emotions and user experience.
@@ 2 The OV-MERD Dataset Construction
Although the concept of OV-MER is intuitive and holds great promise, its practical implementation faces significant challenges. The main difficulty lies in the broad and subtle range of human emotions, making comprehensive labeling a complex task. Traditional annotation methods are limited by their predefined emotion categories, which are often insufficient for the needs of OV-MER. In Figure 2, we propose a human-LLM collaboration strategy that consists of two steps: CLUE-Multi generation and emotion label extraction. Ultimately, we create a dataset called OV-MERD, which offers a richer set of emotions compared to existing datasets (see Table 1). This dataset is an extension of MER2023 [9]. Specifically, MER2023 is collected from movies and TV series, and we randomly select a portion of data for further annotation to build our OV-MERD.
Table 1: Dataset comparison. See Appendix G for a comprehensive comparison.
@@
OV-MERD (Ours) A,V,T Discrete Emotion 236
@@
2.1 CLUE-Multi Generation
During the annotation process, we observe that human-LLM collaboration yields more detailed descriptions than the human-only strategy (see Section 5). In this section, we provide a detailed overview of our strategy.
Pre-annotation.
Initially, we attempt to annotate visual and acoustic clues directly. However, the descriptions obtained in this way cannot cover all information. Therefore, we explore using other models for pre-annotation. (1) For video, given the strong visual understanding capabilities of GPT-4V (“gpt-4-vision-preview”), we use it as VLLM for pre-annotation. Since GPT-4V only supports image input, we uniformly sample three frames from each video and input them into GPT-4V. We discuss the reasons for sampling three frames in Appendix J. (2) For audio, we use the open-source SALMONN [19] as ALLM for pre-annotation, as GPT-4V does not support audio input.
Manual Check.
As part of our quality assurance procedures, we perform a detailed examination of the pre-annotated results. For visual clues, GPT-4V may generate hallucinated responses, i.e., clues that do not actually exist. Additionally, there are repeated expressions and some temporal association clues are missing. Therefore, we hire annotators to eliminate errors and duplicates, as well as add missing content. For acoustic clues, ALLM struggles to capture emotion-related paralinguistic features. The main reason is that current ALLM mainly focuses on tasks like ASR or audio event detection [19], with less emphasis on paralinguistic information. Hence, we hire multiple annotators to focus on the speaker’s intonation and other emotion-related paralinguistic clues. To reduce subjective bias, we conduct two rounds of manual checks, each involving different annotators. These annotators are experts in affective computing and are familiar with the definitions of emotions. Ultimately, these checked clues can accurately reflect the video content. Appendix K provides the annotation guideline and layout of the annotation platform.
CLUE-Multi Generation.
We leverage the reasoning capabilities of LLM to merge all clues. Specifically, we use GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as the LLM and ask it to merge textual, acoustic, and visual clues. The output of this process is an emotion-related description, denoted as CLUE-Multi (see Figure 2). In Appendix J, we further discuss the details of this merging process and the reasons behind it. Overall, the above annotation pipeline reflects the collaboration between humans and LLMs. In Section 5, we compare human-only annotation with our strategy, observing that the latter can generate richer emotions.
2.2 Ground-truth OV Label Extraction
Label Extraction.
After that, we use the LLM to extract emotion labels from CLUE-Multi. This process relies on GPT-3.5, which we request to identify emotional states based on the provided descriptions without restricting the label space. See Appendix J for more details.
Language Impact.
We further explore the language impact. In Figure 2, we first extract OV labels from English and Chinese descriptions, obtaining $Y_{\text{EE}}$ and $Y_{\text{CC}}$. Then, we translate them into the other language, yielding $Y_{\text{EC}}$ and $Y_{\text{CE}}$. Next, we measure the similarity between different sets and report results in Figure 2. In Appendix M, we detail our metric calculation process. We observe that the labels extracted from different languages exhibit some differences. For example, the similarity score between $Y_{\text{EE}}$ and $Y_{\text{CE}}$ is 0.82, which may be due to the varying definitions of emotions in different languages. To eliminate language influence and achieve consensus labels, we merge the labels extracted from both languages and conduct manual checks. These checked labels are regarded as the ground truth.
@@ 2.3 OV-MERD Dataset
Finally, we construct a dataset called OV-MERD. This dataset is an extension of MER2023 [9], from which we randomly select a portion of samples for further annotation. Table 1 compares OV-MERD with existing datasets. We observe that our OV-MERD dataset contains 236 emotion categories and most samples have 2 to 4 labels, far exceeding those in current datasets. In Appendix H, we observe that OV-MERD encompasses a broader range of emotions, including some that have been rarely discussed in previous research, such as shy, nervous, and grateful.
@@ 3 Evaluation Metric
Defining evaluation metrics for OV-MER presents significant challenges: (1) OV-MER supports predicting emotions of any category. Thus, the model may predict closely related but differently expressed emotions. To provide more reliable evaluation results, we first group the emotions based on their similarities. (2) OV-MER allows for the prediction of an arbitrary number of labels. Thus, traditional evaluation metrics designed for a fixed number of labels may not be applicable. In this section, we propose set-based evaluation metrics specifically tailored for this task.